OpenStudy (anonymous):

Taylor can make one banner out of 1 3/4 yards of fabric. How many banners can he make with 10 1/2 yards of fabric?

OpenStudy (anonymous):

Keep adding 1 3/4 over and over and COUNT how many times you do it. Once you get a number over 10 1/2, then that's how many banners Taylor can make.

OpenStudy (anonymous):

medal and fan

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

You could divide 10 1/2 by 1 3/4 to get your answer. That will tell you how many banners Taylor can make.

OpenStudy (anonymous):

10 1/2=10.5 1 3/4=1.75 10.5/1.75=6 Taylor can make 6 banners

OpenStudy (bookworm14):

\(\huge\color{aqua}{I~agree~with~1289florida!!!}\color{purple}{\ddot\smile}\)